3.The following information is for Eucha Corp. for the first quarter of the current fiscal year: Actual Results Static Budget Unit sales: Product X 15,000 40,000 Product Y 65,000 60,000 Total 80,000 100,000 Contribution margin per unit: Product X $4 $5 Product Y $3 $2 The sales-mix variance for both products together is: a. $51,000 unfavorable. b. $64,000 unfavorahle
Product X $4 $5 Product Y $3 $2 The sales-mix variance for both products together is a. $51,000 unfavorable. b. $64,000 unfavorable. c. $115,000 unfavorable. d. $115,000 favorable. 4 Using the information in question 3 the sales quantity variance for Prodtuct Y is: $40,000 unfavorable. a. b. $40,000 favorable. $24,000 unfavorable. C. d. $24,000 favorable. 5.Using the information in question 3, the amount of the budgeted contribution margin per composite unit is: a. $3.50. b. $3.20. c. $2.5625. d. none of the above
Answer #1

Answer for Q3: Sales mix variance for both product together is $51,000 unfavorable

Answer for Q4: Sales quantity variance for Product Y is $24,000 unfavorable

Answer for Q5: Budgeted contribution margin per composite unit is $3.20

Concept:

Sales Variances can be understood using the following chart and formulas:

Total Sales Sales Price Sales Volume Sales Mix Sales Quantity (Actual SP - Budgeted SP) X Actual units sold (Actual units - Budgeted units) X Budgeted profit per unit Sales Price- Sales Volume- Sales mix = Sales Qty- (Actual qty mix-Standard qty mix) X Budgeted profit per unit

Detailed Solution to the Problem:

Mix 96 Actual qty in budgeted mix Actual Standard qty 32,000 48,000 80,000 Budget Budget Product X Product Y Total 15,000 65,000 80,000 40,000 60,000 100,000 18.8 81.3 100.0 40.0 60.0 100.0 Contribution p.u Product X Product Y Actual Budget 4 $2 Sales Volume variance = (Actual units-Budgeted units) X Budgeted profit per unit X(15,000 40,000) X $5 Y= (65,000-60,000) X $2 Total (125,000) UF 10,000 F (115,000) UF - Sales Mix variance(Actual qty mix - Standard qty mix) X Budgeted profit per unit x = (15000 . 32,000) x S5- Y = (65,000-48,000) X $2 Total (85,000) UF 34,000F (51,000) UF Sales aty variance(Standard qty mix Budgeted qty mix) X Budgeted profit per unit x = (32000-40,000) X $5 Y = (48,000-60,000) X $2 Total (40,000) UF (24,000) UF (64,000) UF Note: Sales Volume variance : Sales Mix Variance + Sales Qty Variance Budgeted contribution margin per composite unit (A)x (B) Contribution p.u Product X Product Y Total Budget Budgeted Mix Contribution $2.00 $1.20 $3.20 $5.00 $2.00 40.00% 60.00%